OpenStudy (anonymous):

Help please! Write an equation in slope-intercept form of the line with the given slope and point: slope = 3 and (1, -2).

OpenStudy (the_fizicx99):

y-y1=m(x-x1)

OpenStudy (the_fizicx99):

Plug in your values. Hint: Slope is M

OpenStudy (anonymous):

Can you walk me through it?

OpenStudy (the_fizicx99):

Yea, y-(-2)=3(x-1)

OpenStudy (the_fizicx99):

Negative and negative = a positive so, y+2=3(x-1) Now you distribute 3(x-1)

OpenStudy (anonymous):

okay. is that it?

OpenStudy (the_fizicx99):

No, what did you get?

OpenStudy (anonymous):

I still dont get it.

OpenStudy (the_fizicx99):

3*x = 3x 3*(-1) = -3 y+2=3x-3 - 2 -2 y=3x-5

OpenStudy (the_fizicx99):

Understand now?

OpenStudy (anonymous):

yes. thank you